OpenStudy (iunefr):

The projectile is fired with an initial velocity 40 m/s at an angle of elevation of 30degrees on a platform, find the x and y components of the velocity at t=0

OpenStudy (sapphiremoon):

There are two simple equations we use to find this out in algebra-based physics. One finds the x component and the other the y with only the input of the initial velocity and the firing angle. Both of which you have! Yay! \[v_{0x} = v_{0} \cos \theta_{0}\] and \[v_{0y} = v_{0} \sin \theta_{0}\] Substitute your values into those equations and solve: \[v_{0x} = 40 \cos 30 = 34.64\] and \[v_{0y} = 40 \sin 30 = 20\] Tada! :)

OpenStudy (irishboy123):



OpenStudy (sapphiremoon):

@IrishBoy123 lol